OpenStudy (anonymous):

Solve the equation for x: 3x + y = 6. A. x = 2 - 1/3 y B. x = 3y - 6 C. x = 6 - 3y D. x = 3

OpenStudy (anonymous):

3x = 6 - Y X = (6-Y)/3 = 2- 1/3 y